OpenStudy (liliakarina):

Organic chem! I did this excercise, but I'm not sure how to explain the following: RC(O)CH2C(O)R' + R3N --> RC(O)CHC(O)R' + R3NH+ The equilibrium position goes to the right, because the proton is donated to R3N, so R3NH+, though I don't really know how or why? pKa value for RC(O)CH2C(O)R' is 9, and R3NH+ is 12. How do I know those are the indicated pKa values whose difference will result in the equilibrium constant, Keq, and why do I have to expess it like= 10^+3 (12-9=3)? Is RC(O)CH2C(O)R' a Bronsted acid? It donates protons. Please help me :D, I'm having a really, really hard time :'(

OpenStudy (aaronq):

\(R_3N\) is a the general form for a tertiary amine base, look at the attachment i drew the mechanism for the reaction. pKa is the negative logarithm of Ka, so when you convert to Ka, you take the inverse logarithm which is \(\sf\large pKa=-log(Ka)\rightarrow Ka=10^{-pKa}\) When you wanna find the K for a reaction of two K's combine then you multiply those values. \(\sf \large K_{final}=K_1*K_2\) For this reaction one of them is actually acting as a base, you have the use the Kb instead of the Ka: \(\sf \large K_b=\dfrac{1}{Ka}\) \(\sf \large K_{final}=K_1*\dfrac{1}{K_2}=10^{-9}*\dfrac{1}{10^{-12}}=10^3=1000\)